OpenStudy (darkigloo):

Which series convergence cannot be determined by direct comparison with a geometric series or p-series?

OpenStudy (darkigloo):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ n2^n } \] \[\sum_{k=1}^{\infty} \frac{ k }{ k!+4^k } \] \[\sum_{n=1}^{\infty} \frac{ \sin^2n }{ n^2 } \] \[\sum_{k=1}^{\infty}\frac{ k ^{1/3} }{ k^2+k} \] \[\sum_{n=3}^{\infty} \frac{ 3n+5 }{ n(n-1)(n-2) }\] \[\sum_{n=2}^{\infty} \frac{ n^2 }{ n^4 -1 }\]

OpenStudy (darkigloo):

how do i do this?

satellite73 (satellite73):

one at a time

satellite73 (satellite73):

\[\sum_{n=1}^{\infty} \frac{ 1 }{ n2^n }\] each term is less than \(\frac{1}{2^n}\) a convergent geometric series right?

satellite73 (satellite73):

\[\sum_{n=1}^{\infty} \frac{ \sin^2n }{ n^2 }\] compare with \(-\frac{1}{n^2}\) and \(\frac{1}{n^2}\) then squeeze

satellite73 (satellite73):

oh nvm numerator is positive, compare only with \(\frac{1}{n^2}\)

satellite73 (satellite73):

i have no idea how you do this one\[\sum_{k=1}^{\infty} \frac{ k }{ k!+4^k }\] although convergence is pretty obvious

OpenStudy (darkigloo):

why is that obvious?

satellite73 (satellite73):

because both \(k!\) and \(4^k\) grow way way way (way) faster than \(k\) so together they are like overkill

satellite73 (satellite73):

problem you asked though is to compare to p series or geometric, series, and i don't think they compare to either

satellite73 (satellite73):

we can check the others if you like, but they all look like you could compare to p series to me

OpenStudy (darkigloo):

how would i tell if the convergence cannot be determined?

satellite73 (satellite73):

like for example, what you would compare this to \[\sum_{k=1}^{\infty}\frac{ k ^{1/3} }{ k^2+k}\]

satellite73 (satellite73):

they are not asking if convergence cannot be determined they are asking if it cannot be determines by comparison to p series or geometric series

satellite73 (satellite73):

so is it clear what you would compare \[\sum_{k=1}^{\infty}\frac{ k ^{1/3} }{ k^2+k}\] to easily?

OpenStudy (darkigloo):

k/k^2 ? :/

satellite73 (satellite73):

hmm no

satellite73 (satellite73):

first off \(\frac{k}{k^2}=\frac{1}{k}\) maybe it is not clear but we can make it so the degree of the numerator is \(\frac{1}{3}\) and the degree of the denomnator is \(2\) so you would compare it to \[\large\frac{k^{\frac{1}{3}}}{k^2}=\frac{1}{k^{2-\frac{1}{3}}}=\frac{1}{k^{\frac{5}{3}}}\]

satellite73 (satellite73):

which is a p series with \(p>1\) so it converges `

satellite73 (satellite73):

similarly this one \[\sum_{n=2}^{\infty} \frac{ n^2 }{ n^4 -1 }\] compare with \[\frac{n^2}{n^4}=\frac{1}{n^2}\]